# CoLLM: Integrating Collaborative Embeddings into Large Language Models for Recommendation

### v2

@@ Abstract
Leveraging Large Language Models as recommenders, referred to as LLMRec, is gaining traction and brings novel dynamics for modeling user preferences, particularly for cold-start users. However, existing LLMRec approaches primarily focus on text semantics and overlook the crucial aspect of incorporating collaborative information from user-item interactions, leading to potentially sub-optimal performance in warm-start scenarios. To ensure superior recommendations across both warm and cold scenarios, we introduce CoLLM, an innovative LLMRec approach that explicitly integrates collaborative information for recommendations. CoLLM treats collaborative information as a distinct modality, directly encoding it from well-established traditional collaborative models, and then tunes a mapping module to align this collaborative information with the LLM’s input text token space for recommendations. By externally integrating traditional models, CoLLM ensures effective collaborative information modeling without modifying the LLM itself, providing the flexibility to adopt diverse collaborative information modeling mechanisms. Extensive experimentation validates that CoLLM adeptly integrates collaborative information into LLMs, resulting in enhanced recommendation performance. Our implementations are available in anonymous Github: https://github.com/zyang1580/CoLLM.
@@ 4.1 Model Architecture
  • When encoding prompts, alongside the LLMs’ tokenization and embedding for encoding textual information, we employ a conventional collaborative model to generate user/item representations that capture collaborative information, and map them into the token embedding space of the LLM.
@@ 5.1.1 Evaluation Metrics
We employ two commonly used metrics for explicit recommendation to assess the performance of studied methods: AUC and UAUC [49]. AUC is the Area under the ROC Curve, quantifying the overall prediction accuracy. UAUC is derived by first computing the AUC individually for each user over the exposed items and then averaging these results across all users. AUC evaluates the overall ranking quality, and UAUC provides insights into user-level ranking quality.
TABLE II: Overall performance comparison. “Collab.” denotes collaborative methods. “Rel. Imp.” denotes the relative improvement of CoLLM compared to baselines, averaged over the two metrics. For a collaborative method, the Rel. Imp. is computed using CoLLM implemented on it. For LLMRec methods, it is determined by comparing them to CoLLM-MF.
@@
Methods AUC UAUC Rel. Imp. AUC UAUC Rel. Imp.
Collab. MF 0.6482 0.6361 10.3% 0.7134 0.5565 12.8%
LightGCN 0.5959 0.6499 13.2% 0.7103 0.5639 10.7%
SASRec 0.7078 0.6884 1.9% 0.6887 0.5714 8.4%
DIN 0.7166 0.6459 3.1% 0.8163 0.6145 3.2%
LM+Collab. CTRL (DIN) 0.7159 0.6492 2.9% 0.8202 0.5996 4.2%
LLMRec ICL 0.5320 0.5268 33.8% 0.4820 0.4856 48.2%
Prompt4NR (Vicuna) 0.7071 0.6739 2.7% 0.7224 0.5881 10.4%
TALLRec 0.7097 0.6818 1.8% 0.7375 0.5983 8.2%
Ours CoLLM-MF 0.7295 0.6875 - 0.8109 0.6225 -
CoLLM-LightGCN 0.7100 0.6967 - 0.7978 0.6149 -
CoLLM-SASRec 0.7235 0.6990 - 0.7746 0.5962 -
CoLLM-DIN 0.7243 0.6897 - 0.8245 0.6474 -
@@ 5.2.1 Overall Performance
The overall performance comparison between CoallRec and baselines is summarized in Table II. From the table, we have the following observations:
  • When compared to all other methods, the best version of CoLLM outperforms them in both metrics on the two datasets. The results confirm the excellence of our approach.
  • In comparison between the best LLMRec baseline (TALLRec) and collaborative models, most of the time, it outperforms MF, LightGCN, and SASRec but falls short of beating DIN. However, after introducing collaborative information into LLMRec using CoLLM, LLM consistently achieves performance improvements (except for UAUC for SASRec on Amazon-Book) and surpasses the corresponding collaborative baselines. This demonstrates the necessity of incorporating collaborative information.
  • When focusing on LLMRec methods, the ICL method consistently produces the weakest results, in line with previous findings [12]. This underscores LLM’s inherent limitation in recommendation and emphasizes the importance of fine-tuning LLM for recommendation tasks. Interestingly, when we scrutinize Prompt4NR (Vicuna), it not only fine-tunes LLM itself but also incorporates some adaptable prompts, bearing resemblances to CoLLM. Nevertheless, it not only falls short of matching CoLLM’s performance but even lags behind TALLRec, which exclusively fine-tunes LLM. This suggests that the improvements in CoLLM stem from its collaborative information modeling mechanism rather than only relying on adaptively updatable prompts.
@@ 5.3.3 Efficiency
Concerning inference, the cost associated with the CIE module is anticipated to be negligible due to its significantly smaller scale compared to the LLM. The primary additional cost in our methodology arises from the extra tokens required to describe collaborative information in our prompt template (as indicated in the underlined part) in Section 4.1.1. Fortunately, these additional tokens constitute only a small proportion of the total prompts. Consequently, the supplementary inference cost is expected to be relatively modest. As demonstrated in Table VI, CoLLM incurs only a 12.5% increase in total inference cost on average across the two datasets.
TABLE VI: Training and (Total) Inference Time Comparison: TALLRec vs. CoLLM-MF. ”$\Delta$” indicates the relative cost improvement of CoLLM-MF over TALLRec. ”Book” is short for ”Amazon-book”.

### v3

@@ Abstract
Leveraging Large Language Models as recommenders, referred to as LLMRec, is gaining traction and brings novel dynamics for modeling user preferences, particularly for cold-start users. However, existing LLMRec approaches primarily focus on text semantics and overlook the crucial aspect of incorporating collaborative information from user-item interactions, leading to potentially sub-optimal performance in warm-start scenarios. To ensure superior recommendations across both warm and cold scenarios, we introduce CoLLM, an innovative LLMRec approach that explicitly integrates collaborative information for recommendations. CoLLM treats collaborative information as a distinct modality, directly encoding it from well-established traditional collaborative models, and then tunes a mapping module to align this collaborative information with the LLM’s input text token space for recommendations. By externally integrating traditional models, CoLLM ensures effective collaborative information modeling without modifying the LLM itself, providing the flexibility to adopt diverse collaborative information modeling mechanisms. Extensive experimentation validates that CoLLM adeptly integrates collaborative information into LLMs, resulting in enhanced recommendation performance. Our implementations are available in Github: https://github.com/zyang1580/CoLLM.
@@ 4.1 Model Architecture
  • When encoding prompts, alongside the LLMs’ tokenization and embedding look-up for encoding textual information, we employ a conventional collaborative model to generate user/item representations that capture collaborative information, and map them into the token embedding space of the LLM.
@@ 5.1.1 Evaluation Metrics
We employ two commonly used metrics for explicit recommendation to assess the performance of studied methods: AUC, UAUC [49] and NDCG [50]. AUC is the area under the ROC curve that quantifyes the overall prediction accuracy. UAUC is derived by first computing the AUC individually for each user over the exposed items and then averaging these results across all users. NDCG denotes the Normalized Discounted Cumulative Gain metrics. AUC evaluates the overall ranking quality. UAUC and NDCG provide insights into user-level ranking quality.
TABLE II: Overall performance comparison. “Collab.” denotes collaborative methods. “Rel. Imp.” denotes the relative improvement of CoLLM compared to baselines, averaged over the AUC and UAUC metrics. For a collaborative method ”X”, the Rel. Imp. is computed using CoLLM-X; For LLMRec methods, it is determined by comparing them to CoLLM-MF.
@@
Methods AUC UAUC NDCG Rel. Imp. AUC UAUC NDCG Rel. Imp.
Collab. MF 0.6482 0.6361 0.8447 10.3% 0.7134 0.5565 0.8194 12.8%
LightGCN 0.5959 0.6499 0.8564 13.2% 0.7103 0.5639 0.8245 11.0%
SASRec 0.7078 0.6884 0.8612 1.9% 0.6887 0.5714 0.8244 8.4%
DIN 0.7166 0.6459 0.8496 4.9% 0.8163 0.6145 0.8419 3.2%
LM+Collab. CTRL (DIN) 0.7159 0.6492 0.8559 4.6% 0.8202 0.5996 0.8363 4.2%
LLMRec ICL 0.5320 0.5268 0.8114 33.8% 0.4820 0.4856 0.7917 48.2%
Prompt4NR (Vicuna) 0.7071 0.6739 0.8663 2.7% 0.7224 0.5881 0.8346 10.4%
TALLRec 0.7097 0.6818 0.8711 1.8% 0.7375 0.5983 0.8361 8.2%
Ours CoLLM-MF 0.7295 0.6875 0.8714 - 0.8109 0.6225 0.8457 -
CoLLM-LightGCN 0.7100 0.6967 0.8740 - 0.8026 0.6149 0.8411 -
CoLLM-SASRec 0.7235 0.6990 0.8765 - 0.7746 0.5962 0.8361 -
CoLLM-DIN 0.7353 0.6923 0.8735 - 0.8245 0.6474 0.8550 -
@@ 5.2.1 Overall Performance
The overall performance comparison between CoLLM and baselines is summarized in Table II (where the CoLLM-DIN results on ML-1M are obtained by tuning the whole CIE). From the table, we have the following observations:
  • When compared to baselines, the best version of CoLLM outperforms them in both metrics on the two datasets. The results confirm the excellence of our approach.
  • In comparison between the best LLMRec baseline (TALLRec) and collaborative models, it outperforms MF, LightGCN, and SASRec but falls short of beating DIN. However, after introducing collaborative information into LLMRec using CoLLM, LLM consistently achieves performance improvements (except for UAUC/NDCG for SASRec on Amazon-Book) and surpasses the corresponding collaborative baselines. This demonstrates the necessity of incorporating collaborative information.
  • When focusing on LLMRec methods, the ICL method consistently produces the weakest results, in line with previous findings [12]. This underscores LLM’s inherent limitation in recommendation and emphasizes the importance of tuning LLM for recommendation tasks. Interestingly, when we scrutinize Prompt4NR (Vicuna), it not only fine-tunes LLM itself but also incorporates some adaptable prompts, bearing resemblances to CoLLM. Nevertheless, it not only falls short of matching CoLLM’s performance but even lags behind TALLRec, which exclusively fine-tunes LLM. This suggests that the improvements in CoLLM stem from its collaborative information modeling mechanism rather than only relying on adaptively updatable prompts.
@@ 5.3.3 Efficiency
Concerning inference, the cost associated with the CIE module is anticipated to be negligible due to its significantly smaller scale compared to the LLM. The primary additional cost in our method arises from the extra tokens required to describe collaborative information in our prompt template (as indicated in the underlined part) in Section 4.1.1. These additional tokens constitute only a small proportion of the total prompts. Consequently, the supplementary inference cost is expected to be relatively modest. As demonstrated in Table VI, CoLLM incurs only a 12.5% increase in total inference cost on average across the two datasets.
TABLE VI: Training and Total Inference Time Comparison: TALLRec vs. CoLLM-MF. ”$\Delta$” indicates the relative cost improvement of CoLLM-MF over TALLRec. ”Book” is short for ”Amazon-book”.